a compound contain 90% of carbon and 10%
hydrogen.then  what will be the empirical formula of compound

Grade:10

2 Answers

Harishwar IIT Roorkee
9 years ago
Hi Student,
Given C 90%, H10%.
So Moles of element C = 90/12 =7.5
So Moles of element H = 10/1 =10

Simplest molar ration of carbon= 7.5/7.5=1

Simplest molar ration of H= 10/7.5=1.33

whole no ratio element C= 1*3=3

whole no ratio element H= 1.33*3=4

so its C3H4
Ravi
9 years ago
Convert it to fraction and consider the % weight to be the share of weight in the compound by individual elements. DIvide the fractions with their atomic weight. Simplify the resultants of both the elements and rewrite them in simplest forms, after comparing with each other and then, approximating(only if needed) them to the nearest integer. These for both the elements will be the number of atoms in the compound. Write them as the formula (Symbol followed by the number obtained for each element in subscript) to get the emperical formula.